Seanna​ O'Brien receives pension payments of ​$1,610 at the end of every three months from a retirement fund of ​$52,000. The fund earns 11​% compounded quarterly.

​(a) How many payments will Seanna​ receive?

​(b) What is the size of the final pension​ payment?
Answer #1

Qa:

This is an annuity problem and where Present value of annuity=52000

PV of annuity= A[1-(1+r)^-n]/r

where A=annuity payment ;r= interest rate quarterly= 11/4=2.75%

n= no of quarters=?

52000=1610*[1-(1+2.75%)^-n]/.0275

By equating the values or using trial and error;

n= 80.8 = 81 payments (appx)

Qb:

Size of the final pension payment; we need to find future value of annuity payement

FV of annuity =A*[(1+r)^n-1]/r

=1610*[(1+2.75%)^81 -1]/.0275

=468467.70

----

PV of annuity taking n=81; PV=52041.68

ie; in last term only 1610-41.68 =1568.32 has to be paid.